OpenStudy (anonymous):

three-fourths of a number is 8. what is the number?

OpenStudy (anonymous):

let the number be 'x' \[(3/4)x=8\] now solve for x

OpenStudy (sasogeek):

32/3=x

OpenStudy (anonymous):

you can use a cross-multiply rule here, so (3/4)x = 8 3x = 32 x = 32/3

OpenStudy (anonymous):

how do you solve for x?

OpenStudy (chaise):

When solving for unknowns in the equations we do the inverse operation. For example: 50x = 100 Would you agree if we added 50 to both sides of the equations, the equation would be equal? 50 + 50x = 150 What if we take back away the 50? 50x = 100 What if we divide by 50 from both sides? 50x/50 = 100/50 we find that x = 2 Does this make sense? This is how we solve for unknowns in linear equations, with one unknown.

myininaya (myininaya):

\[\frac{3}{4}x=8 \] to get x by itself we have to look to see what we have to undo notice x is being divided by 4 so multiply both sides by 4 \[4 \cdot \frac{3}{4}x=8 \cdot 4\] 3x=32 now notice x is being multiplied 3 4 so we must divide by 3 \[\frac{\not{3}x}{\not{3}}=\frac{32}{3}\]

myininaya (myininaya):

that is suppose to say: now notice x is being multiplied 3 so we must divide by 3